A lawnmower engine with an efficiency of 0.26 rejects 4700 J of heat every second. What is the magnitude of the work that the engine does in one second?
Answer #1

Efficiency of the lawnmower engine = \eta = 0.26

Heat rejected by the lawnmower engine per second = Qc = 4700 J

Work done by the engine in one second = W

Heat input to the lawnmower engine per second = Qh

Qh = W + Qc

TA n=

W n = W +Qc

W 0.26 = W + 4700

(0.26)(W + 4700) = W

0.26W + 1222 = W

0.74W = 1222

W = 1651.35 J

Work done by the engine in one second = 1651.35 J